OpenStudy (anonymous):

Find the common difference for the arithmetic sequence -18, -14, -10, -6, -2.

OpenStudy (anonymous):

4

OpenStudy (sasogeek):

subtract the second term from the first term, you have ur common difference

OpenStudy (anonymous):

should it b -4 or 4 just wondering

OpenStudy (anonymous):

-14-(-18) = 4

OpenStudy (sasogeek):

it's the second from the first, not the other way round jimmy

OpenStudy (anonymous):

thats what i've done

OpenStudy (sasogeek):

second from first means first comes first. 2 from 4 means 4-2 :)

OpenStudy (anonymous):

no sorry second - first

OpenStudy (sasogeek):

i think common difference is an absolute value

OpenStudy (anonymous):

o i thought it meant what do u subtract to get the next value. im not sure tho

OpenStudy (sasogeek):

it's just the distance between any two consecutive pregressions

OpenStudy (anonymous):

ok with that answer is 4

OpenStudy (anonymous):

a, a+ d, a+ 2d ,a+ 3d d = (a+d) - a

OpenStudy (anonymous):

yes - the answer is +4

OpenStudy (anonymous):

or if u like 4!